OpenStudy (anonymous):

Solve the system using substitution. What is the x- or y-coordinate of the solution? -2x - y = 3 3x + y = 2 A. x = 5 B. x = 7 C. x = 13/2 D. x = 15/2

ganeshie8 (ganeshie8):

Add both the equations vertically : ``` -2x - y = 3 3x + y = 2 --------------- ```

OpenStudy (anonymous):

A?

ganeshie8 (ganeshie8):

Yes! may i knw how you got A ?

OpenStudy (anonymous):

because 3 + 2 = 5

ganeshie8 (ganeshie8):

Excellent!